# On the Use of Hierarchical Vision Foundation Models for Low-Cost Human Mesh Recovery and Pose Estimation

### v1

@@ 3.1 ViTPose [102]
Smaller ViTPose-based encoders can be integrated into the HMR2.0 framework in a straightforward manner. Based on the encoder and decoder parameter sizes of the original HMR2.0 with the ViTPose-H encoder, we construct three additional HMR2.0 variants i.e., HMR2.0-L, HMR2.0-B and HMR2.0-S. Their architectural specifications are summarized in Table 1. Each model is initialized with pretrained weights from the corresponding ViTPose encoder and trained under the same settings as in [26].
@@ 4 Hierarchical VFM as HMR / HPE Encoder
Following this principle, we instantiate HMR and HPE models based on the HMR2.0 and ViTPose frameworks using three hierarchical VFMs: Swin [56], GMF [24], and VM [55]. For clarity, we adopt a consistent naming convention where each model name comprises the encoder identifier, followed by the task identifier (i.e., Pose for HPE or HMR2.0 for HMR), the encoder size (i.e., Base (B), Small (S), or Tiny (T)), and a suffix indicating the number of encoder stages used. For example, SwinPose-S-S3 denotes an HPE model that uses the small variant of Swin Transformer up to Stage 3, while VMHMR2.0-T-S2 refers to an HMR model based on the tiny variant of VMamba using up to Stage 2. For the decoder, we adopt the same architecture as ViTPose or HMR2.0, matched to the corresponding encoder size. Therefore, models such as GMFPose-B-S4, GMFPose-B-S3, and GMFPose-B-S2 share the same decoder architecture, regardless of the number of encoder stages utilized. Note that for HPE, we follow the decoder design of the ViTPose variants [102]. For HMR, the decoder architecture corresponding to each encoder size (i.e., base and small variants of each VFM) is consistent with those defined in Section 3.2.1 (cf. Table 1). Additionally, for tiny encoder models, we use the same decoder architecture as that of the small variants.
@@ 5.1 Setting
We follow the training and evaluation protocols established by our baselines [102, 26] for HMR and HPE. For HPE, we use the COCO dataset [53] for both training and evaluation. All the hierarchical-VFM-based encoders are initialized with ImageNet-1K pre-trained weights provided by their respective official repositories [^3] https://github.com/microsoft/Swin-Transformer [^4] https://github.com/AILab-CVC/GroupMixFormer [^5] https://github.com/MzeroMiko/VMamba. HPE performance is evaluated using Average Precision (AP) and Average Recall (AR) metrics. For HMR, we train our models on a mixed dataset comprising Human3.6M [29], MPI-INF-3DHP [61], COCO [53], MPII [3], InstaVariety [33], AVA [27], and AI Challenger [96]. We evaluate 2D pose estimation accuracy using LSP-Extended [31], COCO-val [53], and PoseTrack-val [4], reporting the Percentage of Correct Keypoints (PCK) of reprojected keypoints at thresholds 0.05 and 0.1. For 3D pose accuracy, we use the 3DPW-test [38] and Human3.6M-val [29] datasets, reporting both Mean Per Joint Position Error (MPJPE) and Procrustes Aligned MPJPE (PA-MPJPE). All models are trained using 8 A100 GPUs. Inference speed, measured in frames per second (FPS), is evaluated using a single A100 GPU.
@@ 5.2 Truncated Hierarchical VFMs
where $\mathcal{D}^{\text{P,2D}}$, $\mathcal{D}^{\text{M,2D}}$, and $\mathcal{D}^{\text{M,3D}}$ denote the datasets for evaluating pose estimation by HPE, 2D pose estimation by HMR, and 3D pose estimation by HMR, respectively [^6] Specifically, in this study $\mathcal{D}^{\text{HPE,2D}}$ consists of COCO-val [53], $\mathcal{D}^{\text{HMR,2D}}$ consists of LSP-Extended [31], COCO-val [53], and PoseTrack-val [4], and $\mathcal{D}^{\text{HMR,3D}}$ consists of 3DPW-test [38] and Human3.6M-val [29]. . Note that higher values of $\Phi^{\text{HPE}}$ and $\Phi^{\text{HMR,2D}}$ indicate better performance, while for $\Phi^{\text{HMR,3D}}$, lower values are preferable.
@@ 5.3 Comparison to Existing Methods
The left panel of Figure 1 illustrates the 3D pose estimation accuracy of HMR models on the Human3.6M dataset [29], including selected hierarchical-VFM-based models, ViT-based HMR2.0 variants, and existing approaches [51, 11, 16]. The right panel presents the 2D pose estimation performance of HPE models, namely, hierarchical-VFM-based models, small ViTPose variants, and existing methods [45, 41, 101], on the COCO dataset [53]. In the left panel, models positioned closer to the bottom-left corner are both more accurate and computationally efficient, while in the right panel, those closer to the top-left corner are preferred. Results in these figures demonstrate that our hierarchical-VFM-based models generally achieve lower PA-MPJPE with smaller model sizes and reduced computational cost for HMR, and offer competitive accuracy for HPE. Despite their simplicity, the findings suggest that hierarchical-VFM-based models can serve as strong and efficient baselines for both HPE and HMR tasks.
@@ 6 Conclusion
ViTPose-H 637.2 125.9 67 79.1 84.1
ViTPose-L 308.5 61.6 126 78.3 83.5
ViTPose-B 90.0 18.5 393 75.8 81.1
ViTPose-S 24.3 5.6 936 73.8 79.2
@@
VMambaPose-B-S4 92.8 16.4 469 75.9 81.0
VMambaPose-B-S3 58.6 14.7 532 76.0 81.1
VMambaPose-B-S2 6.2 4.6 1046 72.1 77.3
VMambaPose-S-S4 53.2 9.7 585 75.2 80.4
VMambaPose-S-S3 33.9 8.8 660 75.2 80.3
VMambaPose-S-S2 4.4 3.1 1234 69.6 74.9
VMambaPose-T-S4 33.3 6.0 951 75.2 80.3
VMambaPose-T-S3 17.0 5.2 1067 75.3 80.4
VMambaPose-T-S2 4.1 2.7 1518 67.9 73.4
Table 3: Pose estimation results of HPE models on the COCO-val dataset [53]. Models highlighted with gray cells are evaluated using publicly available pretrained weights, while the others are trained in our environment.
LSP-Extended [31] COCO [53] PoseTrack [4] 3DPW [38] Human3.6M [29]
@@
ViT-H 670.5 125.6 62 53.3 82.4 86.1 96.2 89.7 97.7 81.3 54.3 49.7 32.1
ViT-L 322.4 60.6 118 50.7 80.3 85.1 95.8 89.5 97.7 80.6 53.5 57.0 35.2
ViT-B 92.8 17.3 366 42.8 74.0 82.1 94.6 85.9 96.2 80.0 52.7 61.3 38.5
ViT-S 24.0 4.5 834 38.4 70.2 80.0 93.5 84.6 95.5 81.9 53.5 63.3 40.5
Swin-B-S4 95.5 18.0 265 45.4 76.5 82.4 94.7 86.2 96.3 82.2 54.6 51.5 34.3
Swin-B-S3 66.1 16.2 289 43.9 75.4 82.6 94.6 86.7 96.2 81.1 54.0 52.1 34.8
Swin-B-S2 9.3 3.1 881 24.8 53.5 67.0 87.2 73.3 90.6 94.2 59.1 69.1 48.3
Swin-S-S4 52.3 10.2 377 43.6 75.7 82.9 94.6 86.9 96.4 81.3 54.4 52.1 34.8
Swin-S-S3 35.8 9.1 408 41.0 73.7 82.0 94.4 86.3 96.0 80.9 53.3 52.6 35.2
Swin-S-S2 3.8 1.7 1075 18.9 44.7 60.4 82.8 67.9 87.6 103.5 63.9 73.3 54.7
Swin-T-S4 31.0 5.2 611 39.7 71.8 80.6 93.8 84.5 95.6 83.5 55.4 52.0 36.1
Swin-T-S4 14.5 4.2 695 36.9 67.8 78.5 92.9 83.1 94.9 83.3 54.4 55.3 38.5
Swin-T-S4 3.8 1.7 1076 20.2 46.3 61.1 83.5 68.6 87.9 100.3 62.0 73.9 54.1
GMF-B-S4 52.4 17.3 160 48.5 79.2 84.2 95.4 88.3 96.9 82.3 54.8 50.3 32.7
GMF-B-S3 29.0 16.1 169 46.9 78.1 83.9 95.2 88.1 96.9 80.4 53.0 53.0 33.8
GMF-B-S2 14.0 13.1 199 41.9 73.8 82.2 94.7 86.7 96.3 79.1 52.3 55.0 35.4
GMF-S-S4 24.8 5.1 425 46.0 77.3 82.7 94.8 86.7 96.3 82.8 54.6 52.1 34.7
GMF-S-S3 19.3 4.8 453 44.2 75.2 82.4 94.7 87.0 96.2 81.4 54.1 53.2 35.4
GMF-S-S2 4.2 1.8 772 31.8 62.0 75.1 91.3 80.6 93.6 85.6 55.0 59.4 41.8
GMF-T-S4 13.1 3.7 380 44.6 76.1 82.5 94.7 86.6 96.2 82.1 54.7 52.5 35.1
GMF-T-S3 10.1 3.5 417 41.7 73.4 81.9 94.3 86.6 96.1 80.3 53.6 54.1 36.4
GMF-T-S2 4.1 2.3 559 32.4 62.9 75.5 91.5 81.2 93.9 83.3 54.0 60.3 41.3
VMamba-B-S4 96.3 15.3 433 45.9 77.7 83.7 95.2 88.1 96.8 81.5 54.3 48.5 34.0
VMamba-B-S3 62.1 13.6 487 44.6 76.1 82.9 95.0 87.0 96.5 80.7 53.6 51.6 34.0
VMamba-B-S2 9.8 3.5 943 32.8 63.0 75.5 91.5 81.1 93.4 83.3 54.2 57.0 40.0
VMamba-S-S4 52.9 8.6 540 46.1 77.1 83.4 94.9 87.8 96.6 81.4 54.3 53.0 34.8
VMamba-S-S3 33.6 7.7 613 44.7 75.6 82.6 94.9 86.5 96.4 81.2 54.1 53.2 35.8
VMamba-S-S2 4.1 2.0 1146 27.3 56.1 71.0 89.2 77.0 92.0 87.1 55.9 60.8 45.2
VMamba-T-S4 33.0 4.9 751 43.5 74.6 82.5 94.5 86.6 96.3 82.0 54.1 51.8 34.8
VMamba-T-S3 14.5 4.2 856 41.8 72.6 81.1 94.2 85.6 95.9 80.3 53.1 53.8 34.9
VMamba-T-S2 3.8 1.6 1402 25.1 53.3 69.7 88.3 76.9 91.6 88.3 57.1 63.7 47.2
Table 4: 2D and 3D pose estimation results of HMR models across five datasets. In this table, “P” denotes PCK, “M” denotes MPJPE, and “PA-M” denotes PA-MPJPE. Models highlighted with gray cells are evaluated using publicly available pretrained weights, while the others are trained in our environment.

### v2

@@ 3.1 ViTPose [xu+2022neurips]
Smaller ViTPose-based encoders can be integrated into the HMR2.0 framework in a straightforward manner. Based on the encoder and decoder parameter sizes of the original HMR2.0 with the ViTPose-H encoder, we construct three additional HMR2.0 variants i.e., HMR2.0-L, HMR2.0-B and HMR2.0-S. Their architectural specifications are summarized in Table 1. Each model is initialized with pretrained weights from the corresponding ViTPose encoder and trained under the same settings as in [goel+2023iccv].
@@ 4 Hierarchical VFM as HMR / HPE Encoder
Following this principle, we instantiate HMR and HPE models based on the HMR2.0 and ViTPose frameworks using three hierarchical VFMs: Swin [liu+2021iccv], GMF [ge+2023arxiv], and VM [liu+2024neurips]. For clarity, we adopt a consistent naming convention where each model name comprises the encoder identifier, followed by the task identifier (i.e., Pose for HPE or HMR2.0 for HMR), the encoder size (i.e., Base (B), Small (S), or Tiny (T)), and a suffix indicating the number of encoder stages used. For example, SwinPose-S-S3 denotes an HPE model that uses the small variant of Swin Transformer up to Stage 3, while VMHMR2.0-T-S2 refers to an HMR model based on the tiny variant of VMamba using up to Stage 2. For the decoder, we adopt the same architecture as ViTPose or HMR2.0, matched to the corresponding encoder size. Therefore, models such as GMFPose-B-S4, GMFPose-B-S3, and GMFPose-B-S2 share the same decoder architecture, regardless of the number of encoder stages utilized. Note that for HPE, we follow the decoder design of the ViTPose variants [xu+2022neurips]. For HMR, the decoder architecture corresponding to each encoder size (i.e., base and small variants of each VFM) is consistent with those defined in Section 3.2.1 (cf. Table 1). Additionally, for tiny encoder models, we use the same decoder architecture as that of the small variants.
@@ 5.1 Setting
We follow the training and evaluation protocols established by our baselines [xu+2022neurips, goel+2023iccv] for HMR and HPE. For HPE, we use the COCO dataset [lin+2014eccv] for both training and evaluation. All the hierarchical-VFM-based encoders are initialized with ImageNet-1K pre-trained weights provided by their respective official repositories [^3] https://github.com/microsoft/Swin-Transformer [^4] https://github.com/AILab-CVC/GroupMixFormer [^5] https://github.com/MzeroMiko/VMamba. HPE performance is evaluated using Average Precision (AP) and Average Recall (AR) metrics. For HMR, we train our models on a mixed dataset comprising Human3.6M [ionescu+2014tpami], MPI-INF-3DHP [mehta+20173dv], COCO [lin+2014eccv], MPII [andriluka+2014cvpr], InstaVariety [kanazawa+2019cvpr], AVA [gu+2018cvpr], and AI Challenger [wu+2019icme]. We evaluate 2D pose estimation accuracy using LSP-Extended [johnson+2011cvpr], COCO-val [lin+2014eccv], and PoseTrack-val [andriluka+2018cvpr], reporting the Percentage of Correct Keypoints (PCK) of reprojected keypoints at thresholds 0.05 and 0.1. For 3D pose accuracy, we use the 3DPW-test [kolotouros+2019iccv] and Human3.6M-val [ionescu+2014tpami] datasets, reporting both Mean Per Joint Position Error (MPJPE) and Procrustes Aligned MPJPE (PA-MPJPE). All models are trained using 8 A100 GPUs. Inference speed, measured in frames per second (FPS), is evaluated using a single A100 GPU.
@@ 5.2 Truncated Hierarchical VFMs
where $\mathcal{D}^{\text{P,2D}}$, $\mathcal{D}^{\text{M,2D}}$, and $\mathcal{D}^{\text{M,3D}}$ denote the datasets for evaluating pose estimation by HPE, 2D pose estimation by HMR, and 3D pose estimation by HMR, respectively [^6] Specifically, in this study $\mathcal{D}^{\text{HPE,2D}}$ consists of COCO-val [lin+2014eccv], $\mathcal{D}^{\text{HMR,2D}}$ consists of LSP-Extended [johnson+2011cvpr], COCO-val [lin+2014eccv], and PoseTrack-val [andriluka+2018cvpr], and $\mathcal{D}^{\text{HMR,3D}}$ consists of 3DPW-test [kolotouros+2019iccv] and Human3.6M-val [ionescu+2014tpami]. . Note that higher values of $\Phi^{\text{HPE}}$ and $\Phi^{\text{HMR,2D}}$ indicate better performance, while for $\Phi^{\text{HMR,3D}}$, lower values are preferable.
@@ 5.3 Comparison to Existing Methods
The left panel of Figure 1 illustrates the 3D pose estimation accuracy of HMR models on the Human3.6M dataset [ionescu+2014tpami], including selected hierarchical-VFM-based models, ViT-based HMR2.0 variants, and existing approaches [lin+2021cvpr, cho+2022eccv, dou+2023iccv]. The right panel presents the 2D pose estimation performance of HPE models, namely, hierarchical-VFM-based models, small ViTPose variants, and existing methods [li+2022eccv, li+2022sensors, xu+2021cvpr], on the COCO dataset [lin+2014eccv]. In the left panel, models positioned closer to the bottom-left corner are both more accurate and computationally efficient, while in the right panel, those closer to the top-left corner are preferred. Results in these figures demonstrate that our hierarchical-VFM-based models generally achieve lower PA-MPJPE with smaller model sizes and reduced computational cost for HMR, and offer competitive accuracy for HPE. Despite their simplicity, the findings suggest that hierarchical-VFM-based models can serve as strong and efficient baselines for both HPE and HMR tasks.
@@ 6 Conclusion
ViTPose-H [xu+2022neurips] 637.2 125.9 67 79.1 84.1
ViTPose-L [xu+2022neurips] 308.5 61.6 126 78.3 83.5
ViTPose-B [xu+2022neurips] 90.0 18.5 393 75.8 81.1
ViTPose-S [xu+2022neurips] 24.3 5.6 936 73.8 79.2
@@
VMPose-B-S4 92.8 16.4 469 75.9 81.0
VMPose-B-S3 58.6 14.7 532 76.0 81.1
VMPose-B-S2 6.2 4.6 1046 72.1 77.3
VMPose-S-S4 53.2 9.7 585 75.2 80.4
VMPose-S-S3 33.9 8.8 660 75.2 80.3
VMPose-S-S2 4.4 3.1 1234 69.6 74.9
VMPose-T-S4 33.3 6.0 951 75.2 80.3
VMPose-T-S3 17.0 5.2 1067 75.3 80.4
VMPose-T-S2 4.1 2.7 1518 67.9 73.4
Table 3: Pose estimation results of HPE models on the COCO-val dataset [lin+2014eccv]. Models shaded with gray cells are evaluated using publicly available pretrained weights, while the others are trained in our environment.
LSP-Extended [johnson+2011cvpr] COCO [lin+2014eccv] PoseTrack [andriluka+2018cvpr] 3DPW [kolotouros+2019iccv] Human3.6M [ionescu+2014tpami]
@@
HMR2.0 [goel+2023iccv] 670.5 125.6 62 53.3 82.4 86.1 96.2 89.7 97.7 81.3 54.3 49.7 32.1
HMR2.0-L 322.4 60.6 118 50.7 80.3 85.1 95.8 89.5 97.7 80.6 53.5 57.0 35.2
HMR2.0-B 92.8 17.3 366 42.8 74.0 82.1 94.6 85.9 96.2 80.0 52.7 61.3 38.5
HMR2.0-S 24.0 4.5 834 39.7 70.8 80.7 93.8 85.5 95.9 81.2 53.4 66.1 41.3
SwinHMR2.0-B-S4 95.5 18.0 265 45.4 76.5 82.4 94.7 86.2 96.3 82.2 54.6 51.5 34.3
SwinHMR2.0-B-S3 66.1 16.2 289 43.9 75.4 82.6 94.6 86.7 96.2 81.1 54.0 52.1 34.8
SwinHMR2.0-B-S2 9.3 3.1 881 24.8 53.5 67.0 87.2 73.3 90.6 94.2 59.1 69.1 48.3
SwinHMR2.0-S-S4 52.3 10.2 377 43.6 75.7 82.9 94.6 86.9 96.4 81.3 54.4 52.1 34.8
SwinHMR2.0-S-S3 35.8 9.1 408 41.0 73.7 82.0 94.4 86.3 96.0 80.9 53.3 52.6 35.2
SwinHMR2.0-S-S2 3.8 1.7 1075 18.9 44.7 60.4 82.8 67.9 87.6 103.5 63.9 73.3 54.7
SwinHMR2.0-T-S4 31.0 5.2 611 39.7 71.8 80.6 93.8 84.5 95.6 83.5 55.4 52.0 36.1
SwinHMR2.0-T-S4 14.5 4.2 695 36.9 67.8 78.5 92.9 83.1 94.9 83.3 54.4 55.3 38.5
SwinHMR2.0-T-S4 3.8 1.7 1076 20.2 46.3 61.1 83.5 68.6 87.9 100.3 62.0 73.9 54.1
GMFHMR2.0-B-S4 52.4 17.3 160 48.5 79.2 84.2 95.4 88.3 96.9 82.3 54.8 50.3 32.7
GMFHMR2.0-B-S3 29.0 16.1 169 46.9 78.1 83.9 95.2 88.1 96.9 80.4 53.0 53.0 33.8
GMFHMR2.0-B-S2 14.0 13.1 199 41.9 73.8 82.2 94.7 86.7 96.3 79.1 52.3 55.0 35.4
GMFHMR2.0-S-S4 24.8 5.1 425 46.0 77.3 82.7 94.8 86.7 96.3 82.8 54.6 52.1 34.7
GMFHMR2.0-S-S3 19.3 4.8 453 44.2 75.2 82.4 94.7 87.0 96.2 81.4 54.1 53.2 35.4
GMFHMR2.0-S-S2 4.2 1.8 772 31.8 62.0 75.1 91.3 80.6 93.6 85.6 55.0 59.4 41.8
GMFHMR2.0-T-S4 13.1 3.7 380 44.6 76.1 82.5 94.7 86.6 96.2 82.1 54.7 52.5 35.1
GMFHMR2.0-T-S3 10.1 3.5 417 41.7 73.4 81.9 94.3 86.6 96.1 80.3 53.6 54.1 36.4
GMFHMR2.0-T-S2 4.1 2.3 559 32.4 62.9 75.5 91.5 81.2 93.9 83.3 54.0 60.3 41.3
VMHMR2.0-B-S4 96.3 15.3 433 45.9 77.7 83.7 95.2 88.1 96.8 81.5 54.3 48.5 34.0
VMHMR2.0-B-S3 62.1 13.6 487 44.6 76.1 82.9 95.0 87.0 96.5 80.7 53.6 51.6 34.0
VMHMR2.0-B-S2 9.8 3.5 943 32.8 63.0 75.5 91.5 81.1 93.4 83.3 54.2 57.0 40.0
VMHMR2.0-S-S4 52.9 8.6 540 46.1 77.1 83.4 94.9 87.8 96.6 81.4 54.3 53.0 34.8
VMHMR2.0-S-S3 33.6 7.7 613 44.7 75.6 82.6 94.9 86.5 96.4 81.2 54.1 53.2 35.8
VMHMR2.0-S-S2 4.1 2.0 1146 27.3 56.1 71.0 89.2 77.0 92.0 87.1 55.9 60.8 45.2
VMHMR2.0-T-S4 33.0 4.9 751 43.5 74.6 82.5 94.5 86.6 96.3 82.0 54.1 51.8 34.8
VMHMR2.0-T-S3 14.5 4.2 856 41.8 72.6 81.1 94.2 85.6 95.9 80.3 53.1 53.8 34.9
VMHMR2.0-T-S2 3.8 1.6 1402 25.1 53.3 69.7 88.3 76.9 91.6 88.3 57.1 63.7 47.2
Table 4: 2D and 3D pose estimation results of HMR models across five datasets. In this table, “P” denotes PCK, “M” denotes MPJPE, and “PA-M” denotes PA-MPJPE. Models shaded with gray cells are evaluated using publicly available pretrained weights, while the others are trained in our environment.